Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Marco Morsiani at Wed Feb 20 22:53:29 MET 2002) Success

Torun (Kazimierz Borkowski at Thu Feb 21 10:24:16 MET 2002) Success
No known problems (the wavelength entry in the feedback do not apply).

Effelsberg (A. Kraus at Thu Feb 21 10:50:47 MET 2002) Minor failures
We had to interrupt the observation between UT1355-1426 (thunderstorm) and between UT1830-1910 (snowfall).

Jodrell_Bank (Alastair Gunn at Fri Feb 22 10:11:02 MET 2002) Minor failures
Telescope parked due to high winds from start of experiment until about 18:56 UT. No other known problems.

Cambridge (Alastair Gunn at Fri Feb 22 10:11:29 MET 2002) Failure
Cambridge did not observe due to serious fault with recorder.

Westerbork (Tony Foley at Fri Feb 22 16:12:06 MET 2002) Minor failures
several scans out of HA range. RT2 and RT5 out at start, and when they returned (after repair RT5 and stable phasing RT2) then RTC out. Online control got stuck 11:33 to 11:47 UT, 14:56 to 15:03:20, 16:09 to 16:20 so no data then.

Onsala (John Conway at Tue Feb 26 17:09:12 MET 2002) Success
No known problems